Doing POX-POY noise measurement as a poor man's FPMI for diagnostic purposes. (Notebook in /opt/rtcds/caltech/c1/Git/40m/measurements/LSC/POX-POY/Noise_Budget.ipynb)
The arms were locked individually using POX11 and POY11. The optical gain was estimated to be by looking at the PDH signal of each arm: the slope was computed by taking the negative peak to positive peak counts and assuming that the arm length change between those peaks is lambda/(2*Finesse), where lambda = 1um and the arm finesse is taken to be 450.
Xarm peak-to-peak counts is ~ 850 while Yarm's is ~ 1100. This gives optical gains of 3.8e11 cts/m and 4.95e11 cts/m respectively.
Next, ETMX actuation TF is measured (attachments 1,2) by exciting C1:LSC-ETMX/Y_EXC and measuring at C1:LSC-X/YARM_IN1_DQ and calibrating with the optical gain.
Using these calibrations I plot the POX-POY (attachment 3) and POX+POY (attachment 4) total noise measurements using two methods:
1. Plotting the calibrated IN and OUT channels of XARM-YARM (blue and orange). Those two curves should cross at the UGF (200Hz in this case).
2. Plotting the calibrated XARM-YARM IN channels times 1-OLTF (black).
The UGF bump can be clearly seen above the true noise in those plots.
However, POX+POY OUT channel looks too high for some reason making the crossing frequency between IN and OUT channels to be ~ 300Hz. Not sure what was going on with this.
Next, I will budget this noise with the individual noise contributions.